Determine the Vertical Asymptote

f(x) = 1-x / x^3 -64

f(x) = 7x + 2/ x^2+1


1
Expert's answer
2021-09-12T23:52:40-0400

1.


"f(x)=\dfrac{1-x}{x^3-64}, x\in \R"

"x^3-64\not=0"

"(x-4)(x^2+x+16)\not=0"

"x\not=4"

"\lim\limits_{x\to4^-}f(x)=\lim\limits_{x\to4^-}\dfrac{1-x}{x^3-64}"

"=\lim\limits_{x\to4^-}\dfrac{1-x}{(x-4)(x^2+x+16)}=\infin"


"\lim\limits_{x\to4^+}f(x)=\lim\limits_{x\to4^+}\dfrac{1-x}{x^3-64}"

"=\lim\limits_{x\to4^+}\dfrac{1-x}{(x-4)(x^2+x+16)}=-\infin"



Vertical asympote: "x=4."


2.


"f(x)=\dfrac{7x+2}{x^2+1}, x\in \R"

"x^2+1>0, x\in\R"

There is no vertical asymptote.